# spacing=nonfrench

### v1

@@ Abstract.
Given a genus $g$ smooth Lefschetz fibration $\pi:M\to S^{2}$ with singular locus $\Delta\subseteq S^{2}$, we describe the subgroup $\operatorname{Br}(\pi)$ of the spherical braid group $\operatorname{Mod}(S^{2},\Delta)$ consisting of braids admitting a lift to a fiber-preserving diffeomorphism of $M$. We develop general methods for showing that the index $\operatorname{Mod}(S^{2},\Delta):\operatorname{Br}(\pi)]$ is infinite. As an application of our methods, we prove that $\operatorname{Mod}(S^{2},\Delta):\operatorname{Br}(\pi)]=\infty$ when $g=1$, when $\pi$ is expressible as a self-fiber sum when $g\geq 2$, or when $\pi$ is a holomorphic genus $g=2$ Lefschetz fibration whose vanishing cycles are nonseparating. In the genus $g=1$ case, we relate the subgroup $\operatorname{Br}(\pi)$ to the action of $\operatorname{Mod}(S,\Delta)$ on the $\operatorname{SL}_{2}$-character variety for $S\setminus\Delta$ and provide an alternate proof of the first application via recent work of Lam–Landesman–Litt [LLL23].
@@ 1. Introduction
    Part (1) of Theorem 1.1 can be deduced from Part (2). Explicitly, every genus 11 Lefschetz fibration is expressed as a self-fiber sum of the rational elliptic surface given by the blowup $\mathcal{B}\ell_{\{p_{1},\ldots,p_{9}\}}\mathbb{P}^{2}$ of $\mathbb{P}^{2}$ along the intersection points of two cubics. The case of the rational elliptic surface itself can be addressed by a similar argument to that used in the proof of Part (2).
@@
Using Chakiris’ classification of holomorphic genus 22 Lefschetz fibrations whose vanishing cycles are nonseparating [Cha83], we are also able to prove the following.
@@
Figure 2. Hurwitz orbits of $\phi_{q_{n}}$ for $n=3$ and 44. Vertex 0 is $\phi_{q_{n}}$, and missing edges indicate representations fixed by the corresponding $\sigma_{i}$. When $n=3$ the orbit has size 88, and when $n=4$ it has size $27$. Figure (B) is the center portion of Figure (A).
@@ Acknowledgements
I thank my advisor Benson Farb for his generosity with his time, feedback, and support. Jointly, I thank Benson Farb and Eduard Looijenga for spending time explaining their work to me. I also thank Aaron Landesman and Daniel Litt for helpful conversations concerning their previous work on character varieties. I appreciate Seraphina Lee and Carlos Servan explaining their work on homologically distinct Lefschetz fibrations, which allowed me to simplify the proof of Theorem $1.$ Thank you to Nick Salter and Dan Margalit, whose comments significantly improved the introduction. Finally, I thank the members of the geometry and topology working group for their time and interest in my project, particularly Ethan Pesikoff and Zhong Zhang. The author was partially supported by an NSF Graduate Research Fellowship under NSF grant no. 2140001.
@@ 2. The Hurwitz Action and $\operatorname{Br}(\pi)$
is the character variety for such local systems. By Theorem 2.1, $\operatorname{Br}(\pi)$ has infinite index if and only if the orbit of $\phi_{\pi}]\in Y_{\Delta}(\mathbb{Z})$ under $\operatorname{Mod}(S^{2},\Delta)$ is infinite. Because the natural map $Y_{\Delta}(\mathbb{Z})\to Y_{\Delta}(\mathbb{C})$ is $\operatorname{Mod}(S^{2},\Delta)$-equivariant, it follows immediately that $\operatorname{Br}(\pi)$ has infinite index inside $\operatorname{Mod}(S^{2},\Delta)$ provided that the $\mathbb{C}$-linear local system has infinite mapping class group orbit. Because the monodromy about a loop $\gamma$ surrounding a single puncture is given by the Dehn twist about the vanishing cycle associated to $\gamma$, which has infinite order, we can apply [LLL23, Corollary 1.1.8], which classifies such rank two local systems of finite mapping class group orbit. In particular, all rank two local systems over the sphere with infinite order monodromy about more than 66 punctures have infinite mapping class group orbit. Hence Part (1) of Theorem 1.1 holds, as any genus one Lefschetz fibration has $12d>6$ singular fibers, since $(\operatorname{SL}_{2}\mathbb{Z})^{ab}\cong\mathbb{Z}/12\mathbb{Z}$.
@@ 4. Finite-Index Exceptions
As a consequence, one finds that the abelianizations $\operatorname{Br}({q_{3}})^{ab},\operatorname{Br}({q_{4}})^{ab}$ are free of rank 22 and that $PB_{n}:PB_{n}\cap\operatorname{Br}({q_{n}})]=[B_{n}:\operatorname{Br}({q_{n}})]$ for $n\leq 4$ (here $PB_{n}$ is the pure braid group).
@@ 5. General Criteria
We’$l$ now devote ourselves to extending Theorem 3.1 over arbitrary compact connected oriented base surfaces $B$. We develop multiple criteria guaranteeing that $\operatorname{Mod}_{\pi}=\operatorname{im}(\operatorname{Mod}(\pi)\to\operatorname{Mod}(B,\Delta))$ has infinite index for a Lefschetz fibration $\pi:M\to B$ with singular fibers along $\Delta\subseteq B\setminus\partial B$. To do so succinctly, we require some additional language.
@@
Lemmas 5.3 and 5.2 also yield a new proof of Part (1) of Theorem $1.$
Alternate proof of Part (1) of Theorem $1.$
@@
The proof is identical to the proof above of Part (1) of Theorem $1.$ Three of the fiber sums are used to produce the five vanishing cycles of Lemma 5.3. The remaining two vanishing cycles serve as a marking.

### v2

@@ Abstract.
Given a genus $g$ smooth Lefschetz fibration $\pi:M\to S^{2}$ with singular locus $\Delta\subseteq S^{2}$, we describe the subgroup $\operatorname{Br}(\pi)$ of the spherical braid group $\operatorname{Mod}(S^{2},\Delta)$ consisting of braids admitting a lift to a fiber-preserving diffeomorphism of $M$. We develop general methods for showing that the index $\operatorname{Mod}(S^{2},\Delta):\operatorname{Br}(\pi)]$ is infinite. As an application of our methods, we prove that $\operatorname{Mod}(S^{2},\Delta):\operatorname{Br}(\pi)]=\infty$ when $g=1$, when $\pi$ is expressible as a self-fiber sum when $g\geq 2$, or when $\pi$ is a holomorphic genus $g=2$ Lefschetz fibration whose vanishing cycles are nonseparating. In the genus $g=1$ case, we relate the subgroup $\operatorname{Br}(\pi)$ to the action of $\operatorname{Mod}(S^{2},\Delta)$ on the $\operatorname{SL}_{2}$-character variety for $S^{2}\setminus\Delta$ and provide an alternate proof of the first application via recent work of Lam–Landesman–Litt [LLL23].
@@ 1. Introduction
    Part (1) of Theorem 1.1 can be deduced from Part (2). Explicitly, every genus $1$ Lefschetz fibration is expressed as a self-fiber sum of the rational elliptic surface given by the blowup $\mathcal{B}\ell_{\{p_{1},\ldots,p_{9}\}}\mathbb{P}^{2}$ of $\mathbb{P}^{2}$ along the intersection points of two cubics. The case of the rational elliptic surface itself can be addressed by a similar argument to that used in the proof of Part (2).
@@
Using Chakiris’ classification of holomorphic genus $2$ Lefschetz fibrations whose vanishing cycles are nonseparating [Cha83], we are also able to prove the following.
@@
Figure 2. Hurwitz orbits of $\phi_{q_{n}}$ for $n=3$ and $4$. Vertex 0 is $\phi_{q_{n}}$, and missing edges indicate representations fixed by the corresponding $\sigma_{i}$. When $n=3$ the orbit has size 88, and when $n=4$ it has size $27$. Figure (B) is the center portion of Figure (A).
@@ Acknowledgements
I thank my advisor Benson Farb for his generosity with his time, feedback, and support. Jointly, I thank Benson Farb and Eduard Looijenga for spending time explaining their work to me. I also thank Aaron Landesman and Daniel Litt for helpful conversations concerning their previous work on character varieties. I appreciate Seraphina Lee and Carlos Servan explaining their work on homologically distinct Lefschetz fibrations, which allowed me to simplify the proof of Theorem 1.1. Thank you to Nick Salter and Dan Margalit, whose comments significantly improved the introduction. Finally, I thank the members of the geometry and topology working group for their time and interest in my project, particularly Ethan Pesikoff and Zhong Zhang. The author was partially supported by an NSF Graduate Research Fellowship under NSF grant no. 2140001.
@@ 2. The Hurwitz Action and $\operatorname{Br}(\pi)$
is the character variety for such local systems. By Theorem 2.1, $\operatorname{Br}(\pi)$ has infinite index if and only if the orbit of $\phi_{\pi}]\in Y_{\Delta}(\mathbb{Z})$ under $\operatorname{Mod}(S^{2},\Delta)$ is infinite. Because the natural map $Y_{\Delta}(\mathbb{Z})\to Y_{\Delta}(\mathbb{C})$ is $\operatorname{Mod}(S^{2},\Delta)$-equivariant, it follows immediately that $\operatorname{Br}(\pi)$ has infinite index inside $\operatorname{Mod}(S^{2},\Delta)$ provided that the $\mathbb{C}$-linear local system has infinite mapping class group orbit. Because the monodromy about a loop $\gamma$ surrounding a single puncture is given by the Dehn twist about the vanishing cycle associated to $\gamma$, which has infinite order, we can apply [LLL23, Corollary 1.1.8], which classifies such rank two local systems of finite mapping class group orbit. In particular, all rank two local systems over the sphere with infinite order monodromy about more than $6$ punctures have infinite mapping class group orbit. Hence Part (1) of Theorem 1.1 holds, as any genus one Lefschetz fibration has $12d>6$ singular fibers, since $(\operatorname{SL}_{2}\mathbb{Z})^{ab}\cong\mathbb{Z}/12\mathbb{Z}$.
@@ 4. Finite-Index Exceptions
As a consequence, one finds that the abelianizations $\operatorname{Br}({q_{3}})^{ab},\operatorname{Br}({q_{4}})^{ab}$ are free of rank $2$ and that $PB_{n}:PB_{n}\cap\operatorname{Br}({q_{n}})]=[B_{n}:\operatorname{Br}({q_{n}})]$ for $n\leq 4$ (here $PB_{n}$ is the pure braid group).
@@ 5. General Criteria
We’ll now devote ourselves to extending Theorem 3.1 over arbitrary compact connected oriented base surfaces $B$. We develop multiple criteria guaranteeing that $\operatorname{Mod}_{\pi}=\operatorname{im}(\operatorname{Mod}(\pi)\to\operatorname{Mod}(B,\Delta))$ has infinite index for a Lefschetz fibration $\pi:M\to B$ with singular fibers along $\Delta\subseteq B\setminus\partial B$. To do so succinctly, we require some additional language.
@@
Lemmas 5.3 and 5.2 also yield a new proof of Part (1) of Theorem 1.1.
Alternate proof of Part (1) of Theorem 1.1.
@@
The proof is identical to the proof above of Part (1) of Theorem 1.1. Three of the fiber sums are used to produce the five vanishing cycles of Lemma 5.3. The remaining two vanishing cycles serve as a marking.